OpenStudy (anonymous):

Okay, I know I'm not crazy on this one. @Kainui (Sorry, but the textbook's been full of mistakes so far, and some are obvious, some make me feel like I'm going insane).

OpenStudy (anonymous):

\[ \iint_S \sqrt{x^2+y^2+z^2}\;dx\wedge dy \]For \(S:x^2+y^2=z^2,\;0\le z\le1\)

OpenStudy (anonymous):

Changing to polar, with \(z=r\) yields, pretty clearly: \[ \sqrt 2\iint_{[0,2\pi]\times [0,1]}d\theta\wedge dr \;r^2 \]

OpenStudy (anonymous):

Which is: \[ \frac{2\sqrt{2}}{3}\pi \]Not the textbook's: \[ \pi\sqrt{2} \]

OpenStudy (anonymous):

(The textbook seems to have missed the \(r\) from the change of variables. But I just want to make sure I'm not completely insane)

OpenStudy (anonymous):

i understand

OpenStudy (kainui):

I think you need to bring in an extra r into your integral, but I'm kind of too hungry to think right now. brb.

OpenStudy (anonymous):

hahahah know feeling cant sleep

OpenStudy (anonymous):

An extra \(r\)? We have one from the root, and the second from the change of variables, which yields the \(r^2\). I'm asking because this is the 3rd mistake in the book, at this point. Unless I'm completely missing something.

OpenStudy (anonymous):

3rd mistake in the book *at the intro,* I should add...

OpenStudy (kainui):

I've not been properly introduced to differential forms. Does dx wedge dy just mean dA normal to the surface of the cone?

OpenStudy (anonymous):

You can just treat it as \(dx\,dy\), as it's the area element.

OpenStudy (anonymous):

Or... well, in this case \(d\theta\;dr\).

OpenStudy (kainui):

Yeah, doesn't dxdy=rdrdθ?

OpenStudy (anonymous):

Yep, that's included in the integral above, I just didn't add it explicitly there.

OpenStudy (anonymous):

("there" being in the post two above the quoted)

OpenStudy (kainui):

I am getting the exact same thing as you, however I feel like something is off in how we're considering the problem. Maybe it is just a typo.

OpenStudy (anonymous):

I'm pretty sure it's just a typo (it seems the author forgot(?) to include the \(r\) from the change to simple polar), but it's like the third one in this intro, that's why I asked. Thanks, though.

OpenStudy (kainui):

Yeah, sorry to hear about that book of yours haha.

OpenStudy (anonymous):

Yeah >.> Well, from looking ahead, the proofs are nice, at least.